# Limit Theorems and Quantitative Statistical Stability for the Equilibrium States of Piecewise Partially Hyperbolic Maps

### v1

@@ 3.1. The W-K like norm
Denote by $\mathcal{AB}_{m}$ the set of signed measures $\mu\in\mathcal{SB}(\Sigma)$ such that its associated positive and negative marginal measures, $\pi_{1\ast}\mu^{+}$ and $\pi_{1\ast}\mu^{-},$ are absolutely continuous with respect to $m$, i.e.,
$\mathcal{AB}_{m}=\{\mu\in\mathcal{SB}(\Sigma):\pi_{1\ast}\mu^{+}<<m\ \ \mathnormal{and}\ \ \pi_{1\ast}\mu^{-}<<m\}.$ (18)
Given a probability measure $\mu\in\mathcal{AB}_{m}$ on $\Sigma$, Theorem (c) describes a disintegration $\left(\{\mu_{\gamma}\}_{\gamma},\mu_{x}\right)$ along $\mathcal{F}$ by a family $\{\mu_{\gamma}\}_{\gamma}$ of probability measures on the stable leaves and, since $\mu\in\mathcal{AB}_{m}$, $\mu_{x}$ can be identified with a non negative marginal density $\phi_{1}:M\longrightarrow\mathbb{R}$, defined almost everywhere, with $|\phi_{1}|_{1}=1$. For a general (non normalized) positive measure $\mu\in\mathcal{AB}_{m}$ we can define its disintegration in the same way. In this case, $\mu_{\gamma}$ are still probability measures, $\phi_{1}$ is still defined and $|\phi_{1}|_{1}=\mu(\Sigma)$.
@@
Let $\pi_{2}:\Sigma\longrightarrow K$ be the projection defined by $\pi_{2}(x,y)=y$. Let $\gamma\in\mathcal{F}$, consider $\pi_{\gamma,2}:\gamma\longrightarrow K$, the restriction of the map $\pi_{2}:\Sigma\longrightarrow K$ to the vertical leaf $\gamma$, and the associated pushforward map $\pi_{\gamma,2\ast}$. Given a positive measure $\mu\in\mathcal{AB}_{m}$ and its disintegration along the stable leaves $\mathcal{F}$, $\left(\{\mu_{\gamma}\}_{\gamma},\mu_{x}=\phi_{1}m\right)$, we define the restriction of $\mu$ on $\gamma$ and denote it by $\mu|_{\gamma}$ as the positive measure on $K$ (not on the leaf $\gamma$) defined, for all mensurable set $A\subset K$, as
@@
For a given signed measure $\mu\in\mathcal{AB}_{m}$ and its Jordan decomposition $\mu=\mu^{+}-\mu^{-}$, define the restriction of $\mu$ on $\gamma$ by
@@
Consider a positive Borelean measure $\mu\in\mathcal{AB}$, and a disintegration $\omega=(\{\mu_{\gamma}\}_{\gamma\in M},\phi_{1})$, where $\{\mu_{\gamma}\}_{\gamma\in M}$ is a family of probabilities on $M\times K$ defined $\widehat{\mu}-$a.e. $\gamma\in M$ (where $\widehat{\mu}:=\pi_{1}{{}_{*}}\mu=\phi_{1}m$) and $\phi_{1}:M\longrightarrow\mathbb{R}$ is a non-negative marginal density. Denote by $\Gamma_{\mu}$ the class of equivalent paths associated to $\mu$
@@
$\mathcal{H}^{\zeta}_{m}=\{\mu\in\mathcal{AB}:|\mu|_{\zeta}<\infty\}.$ (21)
@@
3.5. The $\mathcal{L}_{m}^{\infty}$ and $S^{\infty}$ spaces.
@@
Let $\mathcal{L}_{m}^{\infty}\subseteq\mathcal{AB}_{m}(\Sigma)$ be defined as
$\mathcal{L}_{m}^{\infty}=\left\{\mu\in\mathcal{AB}_{m}:\operatorname{ess\ sup}(||\mu^{+}|_{\gamma}-\mu^{-}|_{\gamma}||_{W}<\infty\right\},$
where the essential supremum is taken over $M$ with respect to $m$. Define the function $||\cdot||_{\infty}:\mathcal{L}_{m}^{\infty}\longrightarrow\mathbb{R}$ by
@@
$S^{\infty}=\left\{\mu\in\mathcal{L}^{\infty};\phi_{1}\in H_{\zeta}\right\},$ (23)
and the function, $||\cdot||_{S^{\infty}}:S^{\infty}\longrightarrow\mathbb{R}$, defined by
@@
4. The actions on $\mathcal{H}^{\zeta}_{m}$, $\mathcal{L}^{\infty}_{m}$ and $S^{\infty}$.
@@ 4.1. The operator
In this section, we define the linear operator $\mathop{\mathrm{F}}_{\Phi,h}:\mathcal{AB}_{m}\longrightarrow\mathcal{AB}_{m}$ by selecting appropriate expressions for the disintegration of the measure $\mathop{\mathrm{F}}_{\Phi,h}\mu$. To achieve this, we also consider the transfer operator of $F$, defined by the standard expression $\mathop{\mathrm{F}}_{\ast}\mu(A)=\mu(F^{-1}(A))$ for any measure $\mu$ and any measurable set $A$.
In what follows, the potentials (defined on $\Sigma$) considered belong to the class
$\mathcal{P}=\{\Phi:\Phi:=\varphi\circ\pi_{1},\varphi\in\mathcal{PH}\}.$ (24)
@@
For a given potential $\Phi\in\mathcal{P}$, define $\mathop{\mathrm{F}_{\Phi,h}}:\mathcal{AB}_{m}\longrightarrow\mathcal{AB}_{m}$ (where $h$ is given by (P1.1)) as the linear operator such that for all $\mu\in\mathcal{AB}_{m}$
@@
For every $\mu\in\mathcal{AB}_{m}$, it holds
@@
For every $\mu\in\mathcal{AB}_{m}$, the restriction of the measure $\mathop{\mathrm{\overline{}}{F}}_{\Phi,h}\mu$ to the leaf $\gamma$, $(\mathop{\mathrm{\overline{}}{F}}_{\Phi,h}\mu)|_{\gamma}$, is given by the expression
@@
Since there is no risk of confusion, we will henceforth denote the operators $\mathop{\mathrm{F}}_{\Phi,h}$ and $\mathop{\mathrm{\overline{}}{F}}_{\Phi,h}$ simply as $\mathop{\mathrm{F}}_{\Phi}$ and $\mathop{\mathrm{\overline{}}{F}}_{\Phi}$, respectively. Moreover, for simplicity, the notations $\mathcal{AB}_{m}$, $\mathcal{L}^{\infty}_{m}$ and $\mathcal{H}^{\zeta}_{m}$ will be abbreviated as $\mathcal{AB}$, $\mathcal{L}^{\infty}$ and $\mathcal{H}^{\zeta}$, respectively.
4.2. On the actions of $\mathop{\mathrm{\overline{}}{F}}_{\Phi}$ on $\mathcal{H}^{\zeta}_{m}$, $\mathcal{L}^{\infty}$ and $S^{\infty}$.
4.2.1. On the actions of $\mathop{\mathrm{\overline{}}{F}}_{\Phi}$ on $\mathcal{L}^{\infty}$ and $S^{\infty}$.
@@
Proposition 4.7 (Lasota-Yorke inequality for $S^{\infty}$).
For all $\mu\in S^{\infty}$, it holds
@@ 4.2.2. On the action of $\mathop{\mathrm{\overline{}}{F}}_{\Phi}$ on $\mathcal{H}^{\zeta}_{m}$
Lemmas 4.8 and 4.9 are precisely Lemmas 3.6 and 3.7 in [4], and their proofs are omitted.
@@
For every $\mu\in\mathcal{AB}$ and $m$-a.e. stable leaf $\gamma\in\mathcal{F}^{s}$, it holds
@@
For the next proposition and henceforth, for a given path $\Gamma_{\mu}^{\omega}\in\Gamma_{\mu}$ (associated with the disintegration $\omega=(\{\mu_{\gamma}\}_{\gamma},\phi_{1})$, of $\mu$), unless written otherwise, we consider the particular path $\Gamma_{\mathop{\mathrm{\overline{}}{F}}_{\Phi}\mu}^{\omega}\in\Gamma_{\mathop{\mathrm{\overline{}}{F}}_{\Phi}\mu}$ defined by the corollaries 4.2 and 4.4, by the expression
@@
The next Theorem 4.15 corresponds to Theorem H in [4], and its proof is given in Subsection 3.1.5 of that paper. For this reason, we omit the proof.
@@
If $F$ satisfies (f1), (f2), (f3), (H1) and (H2), then for each potential $\Phi\in\mathcal{P}$ there exists an unique equilibrium state $\mu_{0}$ in $S^{\infty}$.
@@
Theorem 4.16.
Suppose that $F:\Sigma\longrightarrow\Sigma$ satisfies (f1), (f2), (f3), (H1), (H2) and $(\alpha\cdot L)^{\zeta}<1$. For each potential $\Phi\in\mathcal{P}$ consider its unique equilibrium state $\mu_{0}\in S^{\infty}$. Then $\mu_{0}\in\mathcal{H}_{\zeta}$ and
$|\mu_{0}|_{\zeta}\leq\dfrac{D}{1-\beta},$ (36)
where $D$ and $\beta$ are from Proposition 4.11.
@@
Let us consider the following sets of zero average measures in $S^{\infty}$ defined by
$\mathcal{V}^{\infty}=\{\mu\in S^{\infty}:\phi_{1}\in\operatorname{Ker}(\mathop{\mathrm{P}}_{f})\}.$ (37)
Theorem 4.17 (Exponential convergence to the equilibrium in $\mathcal{L}^{\infty}$ and $\mathcal{S}^{\infty}$).
Suppose that $F$ satisfies (f1), (f2), (f3), (H1) and (H2). Then, there exist $D_{3}\in\mathbb{R}$ and $0\leq\beta_{3}<1$ such that for every signed measure $\mu\in\mathcal{V}^{\infty}$, it holds
@@ 5. Exponential Decay of Correlations and Central Limit Theorem
In this section we establish several limit theorems for the dynamical system $(F,\mu_{0})$. In particular, we show that the system exhibits exponential decay of correlations for observables in the space $L^{1}(\mathcal{F}_{0})$, introduced below, and in the Hölder space $\operatorname{H}_{\zeta}(\Sigma)$. Moreover, we prove that $(F,\mu_{0})$ satisfies the Central Limit Theorem for Hölder observables.
5.1. Exponential Decay of Coorelations over the constant fiber functions
@@
We prove that $(F,\mu_{0})$ exhibits exponential decay of correlations for two classes of observables: the first consists of functions that are constant along the fibers, and the second is the space of Hölder continuous functions. To accomplish this, we need the following lemma.
@@
Since the Hölder constant of $\mu_{0}$ is finite (Theorem 4.16), there exists a disintegration $\omega=(\{\mu_{0,\gamma}\}_{\gamma\in M},\phi_{1})$ of $\mu_{0}$ such that $|\Gamma_{\mu_{0}}^{\omega}|_{\zeta}<+\infty$. Moreover, $\phi_{1}\equiv 1$ $m$-a.e.
@@
Thus, by Theorem 4.16, we have
@@
Theorem A.
Suppose that $F:\Sigma\longrightarrow\Sigma$ satisfies (f1), (f2), (f3), (H1) and (H2) and $(\alpha\cdot L)^{\zeta}<1$. Let $\mu_{0}$ be the unique equilibrium state of $\Phi\in\mathcal{P}$ in $S^{\infty}$. There exists a constant $0<\tau_{2}<1$ such that, for every constant fiber function $\psi:\Sigma\to\mathbb{R}$ satisfying $\psi(\cdot,y)\in L^{1}(m)$ for all $y$, and for every $\varphi\in\operatorname{H}_{\zeta}(\Sigma)$, we have

### v2

@@ 3.1. The W-K like norm
Denote by $\mathbf{AB}_{m}$ the set of signed measures $\mu\in\mathcal{SB}(\Sigma)$ such that its associated positive and negative marginal measures, $\pi_{1\ast}\mu^{+}$ and $\pi_{1\ast}\mu^{-},$ are absolutely continuous with respect to $m$, i.e.,
$\mathbf{AB}_{m}=\{\mu\in\mathcal{SB}(\Sigma):\pi_{1\ast}\mu^{+}<<m\ \ \mathnormal{and}\ \ \pi_{1\ast}\mu^{-}<<m\}.$ (18)
Given a probability measure $\mu\in\mathbf{AB}_{m}$ on $\Sigma$, Theorem (c) describes a disintegration $\left(\{\mu_{\gamma}\}_{\gamma},\mu_{x}\right)$ along $\mathcal{F}$ by a family $\{\mu_{\gamma}\}_{\gamma}$ of probability measures on the stable leaves and, since $\mu\in\mathbf{AB}_{m}$, $\mu_{x}$ can be identified with a non negative marginal density $\phi_{1}:M\longrightarrow\mathbb{R}$, defined almost everywhere, with $|\phi_{1}|_{1}=1$. For a general (non normalized) positive measure $\mu\in\mathbf{AB}_{m}$ we can define its disintegration in the same way. In this case, $\mu_{\gamma}$ are still probability measures, $\phi_{1}$ is still defined and $|\phi_{1}|_{1}=\mu(\Sigma)$.
@@
Let $\pi_{2}:\Sigma\longrightarrow K$ be the projection defined by $\pi_{2}(x,y)=y$. Let $\gamma\in\mathcal{F}$, consider $\pi_{\gamma,2}:\gamma\longrightarrow K$, the restriction of the map $\pi_{2}:\Sigma\longrightarrow K$ to the vertical leaf $\gamma$, and the associated pushforward map $\pi_{\gamma,2\ast}$. Given a positive measure $\mu\in\mathbf{AB}_{m}$ and its disintegration along the stable leaves $\mathcal{F}$, $\left(\{\mu_{\gamma}\}_{\gamma},\mu_{x}=\phi_{1}m\right)$, we define the restriction of $\mu$ on $\gamma$ and denote it by $\mu|_{\gamma}$ as the positive measure on $K$ (not on the leaf $\gamma$) defined, for all mensurable set $A\subset K$, as
@@
For a given signed measure $\mu\in\mathbf{AB}_{m}$ and its Jordan decomposition $\mu=\mu^{+}-\mu^{-}$, define the restriction of $\mu$ on $\gamma$ by
@@
Consider a positive Borelean measure $\mu\in\mathbf{AB}$, and a disintegration $\omega=(\{\mu_{\gamma}\}_{\gamma\in M},\phi_{1})$, where $\{\mu_{\gamma}\}_{\gamma\in M}$ is a family of probabilities on $M\times K$ defined $\widehat{\mu}-$a.e. $\gamma\in M$ (where $\widehat{\mu}:=\pi_{1}{{}_{*}}\mu=\phi_{1}m$) and $\phi_{1}:M\longrightarrow\mathbb{R}$ is a non-negative marginal density. Denote by $\Gamma_{\mu}$ the class of equivalent paths associated to $\mu$
@@
$\mathcal{H}^{\zeta}_{m}=\{\mu\in\mathbf{AB}:|\mu|_{\zeta}<\infty\}.$ (21)
@@
3.5. The $\mathbf{L}_{m}^{\infty}$ and $\mathbf{S}^{\infty}$ spaces.
@@
Let $\mathbf{L}_{m}^{\infty}\subseteq\mathbf{AB}_{m}(\Sigma)$ be defined as
$\mathbf{L}_{m}^{\infty}=\left\{\mu\in\mathbf{AB}_{m}:\operatorname{ess\ sup}(||\mu^{+}|_{\gamma}-\mu^{-}|_{\gamma}||_{W}<\infty\right\},$
where the essential supremum is taken over $M$ with respect to $m$. Define the function $||\cdot||_{\infty}:\mathbf{L}_{m}^{\infty}\longrightarrow\mathbb{R}$ by
@@
$\mathbf{S}^{\infty}=\left\{\mu\in\mathbf{L}^{\infty};\phi_{1}\in H_{\zeta}\right\},$ (23)
and the function, $||\cdot||_{\mathbf{S}^{\infty}}:\mathbf{S}^{\infty}\longrightarrow\mathbb{R}$, defined by
@@
4. The actions on $\mathcal{H}^{\zeta}_{m}$, $\mathbf{L}^{\infty}_{m}$ and $\mathbf{S}^{\infty}$.
@@ 4.1. The operator
In this section, we define the linear operator $\mathop{\mathrm{F}}_{\Phi,h}:\mathbf{AB}_{m}\longrightarrow\mathbf{AB}_{m}$ by selecting appropriate expressions for the disintegration of the measure $\mathop{\mathrm{F}}_{\Phi,h}\mu$. To achieve this, we also consider the transfer operator of $F$, defined by the standard expression $\mathop{\mathrm{F}}_{\ast}\mu(A)=\mu(F^{-1}(A))$ for any measure $\mu$ and any measurable set $A$.
In what follows, we define a set of fibre constant potentials defined on $\Sigma$.
Definition 4.1.
Define
$\mathscr{P}_{\Sigma}:=\left\{\Phi\;;\;\Phi=\varphi\circ\pi_{1},\ \varphi\in\mathscr{P}_{M}\right\}.$ (24)
@@
For a given potential $\Phi\in\mathscr{P}_{\Sigma}$, define $\mathop{\mathrm{F}_{\Phi,h}}:\mathbf{AB}_{m}\longrightarrow\mathbf{AB}_{m}$ (where $h$ is given by (P1.1)) as the linear operator such that for all $\mu\in\mathbf{AB}_{m}$
@@
For every $\mu\in\mathbf{AB}_{m}$, it holds
@@
For every $\mu\in\mathbf{AB}_{m}$, the restriction of the measure $\mathop{\mathrm{\overline{}}{F}}_{\Phi,h}\mu$ to the leaf $\gamma$, $(\mathop{\mathrm{\overline{}}{F}}_{\Phi,h}\mu)|_{\gamma}$, is given by the expression
@@
Since there is no risk of confusion, we will henceforth denote the operators $\mathop{\mathrm{F}}_{\Phi,h}$ and $\mathop{\mathrm{\overline{}}{F}}_{\Phi,h}$ simply as $\mathop{\mathrm{F}}_{\Phi}$ and $\mathop{\mathrm{\overline{}}{F}}_{\Phi}$, respectively. Moreover, for simplicity, the notations $\mathbf{AB}_{m}$, $\mathbf{L}^{\infty}_{m}$ and $\mathcal{H}^{\zeta}_{m}$ will be abbreviated as $\mathbf{AB}$, $\mathbf{L}^{\infty}$ and $\mathcal{H}^{\zeta}$, respectively.
4.2. On the actions of $\mathop{\mathrm{\overline{}}{F}}_{\Phi}$ on $\mathcal{H}^{\zeta}_{m}$, $\mathbf{L}^{\infty}$ and $\mathbf{S}^{\infty}$.
4.2.1. On the actions of $\mathop{\mathrm{\overline{}}{F}}_{\Phi}$ on $\mathbf{L}^{\infty}$ and $\mathbf{S}^{\infty}$.
@@
Proposition 4.8 (Lasota-Yorke inequality for $\mathbf{S}^{\infty}$).
For all $\mu\in\mathbf{S}^{\infty}$, it holds
@@ 4.2.2. On the action of $\mathop{\mathrm{\overline{}}{F}}_{\Phi}$ on $\mathcal{H}^{\zeta}_{m}$
Lemmas 4.9 and 4.10 are precisely Lemmas 3.6 and 3.7 in [4], and their proofs are omitted.
@@
For every $\mu\in\mathbf{AB}$ and $m$-a.e. stable leaf $\gamma\in\mathcal{F}^{s}$, it holds
@@
For the next proposition and henceforth, for a given path $\Gamma_{\mu}^{\omega}\in\Gamma_{\mu}$ (associated with the disintegration $\omega=(\{\mu_{\gamma}\}_{\gamma},\phi_{1})$, of $\mu$), unless written otherwise, we consider the particular path $\Gamma_{\mathop{\mathrm{\overline{}}{F}}_{\Phi}\mu}^{\omega}\in\Gamma_{\mathop{\mathrm{\overline{}}{F}}_{\Phi}\mu}$ defined by the corollaries 4.3 and 4.5, by the expression
@@
Theorem 4.16 below corresponds to Theorem H in [4]. We omit its proof here, as it can be found in [4, Subsection 3.1.5]. This result is fundamental, as it characterizes the conditions under which the probability measure $\mu_{0}$ is an equilibrium state.
@@
Assume that $F$ satisfies conditions (f1), (f2), (f3), and (H1). Then, for each potential $\Phi\in\mathscr{P}_{\Sigma}$, there exists a unique $F$-invariant measure $\mu_{0}\in\mathbf{S}^{1}$, which in addition belongs to $\mathbf{S}^{\infty}$. In particular, if $F$ uniformly contracts all fibres, then $\mu_{0}$ is an equilibrium state.
@@
Theorem A.
Suppose that $F:\Sigma\to\Sigma$ satisfies conditions (f1), (f2), (f3), (H1), (H2), and $(\alpha\cdot L)^{\zeta}<1$. For each potential $\Phi\in\mathscr{P}_{\Sigma}$, let $\mu_{0}$ denote the $F$-invariant measure of Theorem 4.16. Then $\mu_{0}\in\mathcal{H}_{\zeta}^{+}$ and
$|\mu_{0}|_{\zeta}\leq\frac{D}{1-\beta},$ (36)
where $D$ and $\beta$ are the constants given in Proposition 4.12.
@@
Let us consider the following sets of zero average measures in $\mathbf{S}^{\infty}$ defined by
$\mathbf{V}^{\infty}=\{\mu\in\mathbf{S}^{\infty}:\phi_{1}\in\operatorname{Ker}(\mathop{\mathrm{P}}_{f})\}.$ (37)
Theorem 4.17 (Exponential convergence to the equilibrium in $\mathbf{L}^{\infty}$ and $\mathcal{S}^{\infty}$).
Suppose that $F$ satisfies (f1), (f2), (f3), (H1) and (H2). Then, there exist $D_{3}\in\mathbb{R}$ and $0\leq\beta_{3}<1$ such that for every signed measure $\mu\in\mathbf{V}^{\infty}$, it holds
@@ 5. Exponential Decay of Correlations and Central Limit Theorem
In this section, we establish several limit theorems for the dynamical system $(F,\mu_{0})$. In particular, we show that the system exhibits exponential decay of correlations for observables in the space $L^{1}(\mathcal{F}_{0})$, introduced below, and in the Hölder space $\operatorname{H}_{\zeta}(\Sigma)$. Moreover, we prove that $(F,\mu_{0})$ satisfies the Central Limit Theorem for Hölder observables.
5.1. Exponential Decay of Correlations over the constant fibre functions
@@
We prove that $(F,\mu_{0})$ exhibits exponential decay of correlations for two classes of observables: the first consists of functions that are constant along the fibres, and the second is the space of Hölder continuous functions. To accomplish this, we need the following lemma.
@@
Since the Hölder constant of $\mu_{0}$ is finite (Theorem A), there exists a disintegration $\omega=(\{\mu_{0,\gamma}\}_{\gamma\in M},\phi_{1})$ of $\mu_{0}$ such that $|\Gamma_{\mu_{0}}^{\omega}|_{\zeta}<+\infty$. Moreover, $\phi_{1}\equiv 1$ $m$-a.e.
@@
Thus, by Theorem A, we have
@@
Theorem B.
Suppose that $F:\Sigma\longrightarrow\Sigma$ satisfies (f1), (f2), (f3), (H1) and (H2) and $(\alpha\cdot L)^{\zeta}<1$. Let $\mu_{0}$ be the unique $F$-invariant probability in $\mathbf{S}^{\infty}$. There exists a constant $0<\tau_{2}<1$ such that, for every constant fibre function $\psi:\Sigma\to\mathbb{R}$ satisfying $\psi(\cdot,y)\in L^{1}(m)$ for all $y$, and for every $\varphi\in\operatorname{H}_{\zeta}(\Sigma)$, we have